1 This is the first section in file include.tex

Here is some text to create a paragraph or two so that we can see if this works or not. It will be interesting to see if the labels work properly. As I create these testing files I realize that I need to add parsing support for \input and \include in the getSection function. Furthermore, after reading the L^{A}TEX book, I see that \include files will all start on a new page. This is not the case for \input files. Here is a reference to the next section ?? in this file, and one to the first subsection in the first include file include1 ??. This is a reference ?? to the first subsubsection in the file include3 that is included by include2.

1.1 This is the first subsection from file include1.tex

Here is some text to create a paragraph or two so that we can see if this works or not. It will be interesting to see if the labels work properly. As I create these testing files I realize that I need to add parsing support for \input and \include in the getSection function. Furthermore, after reading the L^{A}TEX book, I see that \include files will all start on a new page. This is not the case for \input files. Here is a reference to the next subsection ??.

1.2 This is the second subsection from file include1.tex

Here is some text to create a paragraph or two so that we can see if this works or not. It will be interesting to see if the labels work properly. As I create these testing files I realize that I need to add parsing support for \input and \include in the getSection function. Furthermore, after reading the L^{A}TEX book, I see that \include files will all start on a new page. This is not the case for \input files. Here is a reference to the previous subsection ??.

1.3 This is the first subsection in file include2.tex

Here is some text to create a paragraph or two so that we can see if this works or not. It will be interesting to see if the labels work properly. As I create these testing files I realize that I need to add parsing support for \input and \include in the getSection function. Furthermore, after reading the L^{A}TEX book, I see that \include files will all start on a new page. This is not the case for \input files. Here is a reference to the next subsection ??.

1.4 This is the second subsection in file include2.tex

Here is some text to create a paragraph or two so that we can see if this works or not. It will be interesting to see if the labels work properly. As I create these testing files I realize that I need to add parsing support for \input and \include in the getSection function. Furthermore, after reading the L^{A}TEX book, I see that \include files will all start on a new page. This is not the case for \input files. Here is a reference to the previous subsection ??. Here is a reference to the first section in include1.tex ??.

Now we include another file from within an included file

2 Testing how including files within other files

The trick here is to test the various ways that files might be included within another file.

2.0.1 This is the first subsubsection from file include3.tex

Here is some text to create a paragraph or two so that we can see if this works or not. It will be interesting to see if the labels work properly. As I create these testing files I realize that I need to add parsing support for \input and \include in the getSection function. Furthermore, after reading the L^{A}TEX book, I see that \include files will all start on a new page. This is not the case for \input files. Here is a reference to the next subsubsection ??.

2.0.2 This is the second subsubsection from file include3.tex

Here is some text to create a paragraph or two so that we can see if this works or not. It will be interesting to see if the labels work properly. As I create these testing files I realize that I need to add parsing support for \input and \include in the getSection function. Furthermore, after reading the L^{A}TEX book, I see that \include files will all start on a new page. This is not the case for \input files. Here is a reference to the previous subsubsection ??. Here is a reference to the first subsection from include1.tex ??. This is the content following \input{include3}. There should be no pagebreak.

2.0.3 This is the first subsubsection from file include4.tex

Here is some text to create a paragraph or two so that we can see if this works or not. It will be interesting to see if the labels work properly. As I create these testing files I realize that I need to add parsing support for \input and \include in the getSection function. Furthermore, after reading the L^{A}TEX book, I see that \include files will all start on a new page. This is not the case for \input files. Here is a reference to the next subsubsection ?? and one in the next input file ??.

2.0.4 This is the second subsubsection from file include4.tex

Here is some text to create a paragraph or two so that we can see if this works or not. It will be interesting to see if the labels work properly. As I create these testing files I realize that I need to add parsing support for \input and \include in the getSection function. Furthermore, after reading the L^{A}TEX book, I see that \include files will all start on a new page. This is not the case for \input files. Here is a reference to the previous subsubsection ??. Here is a reference to the first subsection from include1.tex ??. This is the content following \input include4 . There should be no pagebreak.

3 This is the second section in file include.tex

Here is some text to create a paragraph or two so that we can see if this works or not. It will be interesting to see if the labels work properly. As I create these testing files I realize that I need to add parsing support for \input and \include in the getSection function. Furthermore, after reading the L^{A}TEX book, I see that \include files will all start on a new page. This is not the case for \input files. Here is a reference to the first subsection in include2.tex ??. Here is a reference to the first subsection in include1.tex ??. And finally one to the first section of this file ??.

3.0.1 This is the first subsubsection from file include5.tex

Here is a reference to the next subsubsection ?? and one in the previous input file ??. Notice that the reference to the next subsubsection is 3.0.5 and not 3.0.1. This shows sections reset subsection counters, but not subsubsection counters. How bizarre!

3.0.2 This is the second subsubsection from file include5.tex

Furthermore, after reading the L^{A}TEX book, I see that \include files will all start on a new page. This is not the case for \input files. Here is a reference to the previous subsubsection ??. Here is a reference to the first subsection from include1.tex ??. This is the content following \input include5.tex. There should be no pagebreak.